OpenStudy (anonymous):

help please Subject: Hypotenuse. @ali2x2 Question 1 of 5. The legs of a right triangle are 10 centimeters and 24 centimeters long. What is the length of the hypotenuse? 22 centimeters 26 centimeters 30 centimeters 34 centimeters

OpenStudy (ali2x2):

ayy

OpenStudy (anonymous):

The legs of a right triangle are 10 centimeters and 24 centimeters long. What is the length of the hypotenuse? 22 centimeters 26 centimeters 30 centimeters 34 centimeters

OpenStudy (ali2x2):

im not sure 34?

OpenStudy (anonymous):

Pyragothean Theorem A^2 + B^2 = C^2

OpenStudy (anonymous):

10^2 + 24^2 = C^2

OpenStudy (anonymous):

100 + 576 = C^2

OpenStudy (anonymous):

C^2 = 676 C is the length of the hypotenuse, so \[\sqrt{676} = C\] C = 26 centimeters

OpenStudy (anonymous):

Option B.